alones (alones):

IS THE DENOMINATOR 18 HEERE?

OpenStudy (benlindquist):

eh

alones (alones):

\[ 9x^2-5x-1=0 \]

OpenStudy (sweetburger):

denominator of what..

alones (alones):

And my answer is \[\frac{ 5+\sqrt{61} }{ 18 },\frac{ 5-\sqrt{61} }{ ? }\]

alones (alones):

Interesting...

OpenStudy (sweetburger):

yea the denominator is 18

alones (alones):

Wait what?

OpenStudy (atrineas):

Yes it is

OpenStudy (benlindquist):

no

OpenStudy (atrineas):

YES

OpenStudy (atrineas):

If you use the quadrtic formula you will get what she has

OpenStudy (atrineas):

quadratic*

alones (alones):

Wait so i'm wrong

alones (alones):

**right~

OpenStudy (atrineas):

Your correct the denominator would be 18

alones (alones):

Tanks

alones (alones):

@dmezzullo confused me >.>